Consider a CNC machine in an automated production line that has a mean time between failure (MTBF) of 45 hours. What should be the mean time to repair (MTTR, in hours) if we want the CNC's availability to be 0.90? Round your answer to the nearest integer if it is not already an integer
Answer #1

MTBF ( mean time between failure) = 45 hours

Availability ( of machine) = 0.90

We have to find MTTR ( mean time to failure)

As formula for availability is

Availability = MTBF/ ( MTBF+ MTTR)

0.90 = 45 / ( 45+ MTTR)

Thus MTTR = (45/ 0.90) - 45

= 50 - 45 = 5

Thus mean time to failure (MTTR) is 5 hours.

.